OpenStudy (anonymous):

What did I do wrong here??? x/2 + 7 = 9 x/2 + 7 - 7 = 9 - 7 x/2 = 1 2 * x/2 = 2 * 1 x = 2 Check: x/2 + 7 = 9 2/2 + 7 = 8?

Parth (parthkohli):

Isn't that correct?

OpenStudy (anonymous):

no cause when I check it 2/2 + 7 = 8 not 9

Parth (parthkohli):

I mean, \(\dfrac{2}{2 } + 7 = 1 + 7 = 8\) which is correct

Parth (parthkohli):

\[\dfrac{2}{2} =1 \]Right?

OpenStudy (anonymous):

Yeah

Parth (parthkohli):

So \(\dfrac{2}{2} + 7\) is the same as \(1 + 7\), right?

OpenStudy (anonymous):

But I am trying to fin x =? The equation has to =9 not 8

Parth (parthkohli):

Uh oh, I read it wrong.

Parth (parthkohli):

You wrote \(9 - 7 = 1\) in one step. :-)

OpenStudy (anonymous):

oh it = 2

Parth (parthkohli):

exactly!

Parth (parthkohli):

All right, you're good to go from here.

OpenStudy (anonymous):

haha wow! Thanks for the help!! :D

Parth (parthkohli):

Nay problem!